OpenStudy (anonymous):

Given u = log(x) and v = log(y). Express log (y^2/^3√x^4) in terms of u and v.

OpenStudy (anonymous):

@obafemich ;)

OpenStudy (anonymous):

:D

OpenStudy (luigi0210):

Is it \(\LARGE log\frac{y^2}{\sqrt[4]{x^3}}\)?

OpenStudy (anonymous):

switch the 4 and 3 in the denominator

OpenStudy (luigi0210):

Whoops. But all you have to do is use log rules.

OpenStudy (anonymous):

that wouldn't be a problem however i am very bad at math

OpenStudy (anonymous):

what exactly is being asked

OpenStudy (luigi0210):

\[\LARGE log\frac{y^2}{\sqrt[3]{x^4}}=log~y^2-log~x^{4/3}\] And it just wants you to substitute

OpenStudy (anonymous):

O_O woooooooooooooooooow im dumb

OpenStudy (luigi0210):

Naw, it does look scary tho xD Now just apply the exponent rule for logs, sub in and you're done

OpenStudy (anonymous):

@Luigi0210 so in terms of showing my answer do i just jot down the domain?

OpenStudy (loser66):

u = log (x) --> x = \(10^u\)--> \(x^{3/4}= 10^{3u/4}\) v = log(y) --> y = \(10^v\)--> \(\large y^2= 10^{2v}\) now take the quotient \(\large \dfrac{y^2}{x^{3/4}}=\dfrac{10^{2v}}{10^{3u/4}}=10^{2v-3u/4}\) That is the process you express x, y in terms of u and v

OpenStudy (loser66):

and then take log both sides to get 2v -3u/4

OpenStudy (anonymous):

@Loser66 yeai get that but what i dont get is how should i express my answer

OpenStudy (loser66):